New home sales this year in the Sacramento area were 14% fewer than last year. If the number of new homes sold this year was 5645, find the number of new homes sold last year. Round to the nearest whole number.
Answer by stanbon(75887)

New home sales this year in the Sacramento area were 14% fewer than last year. If the number of new homes sold this year was 5645, find the number of new homes sold last year. Round to the nearest whole number.
-----
Let x be # sold last year.
---
Equation:
x = 5645-0.14*5645 = 0.86*5645 = 4855
=================
Cheers,
Stan H.
=================